OpenStudy (amistre64):

Given a set X, if the subsets of X with exactly 3 elements is 14 more than the subsets of X with exactly 2 elements; then how many subsets of X have exactly 4 elements?

OpenStudy (amistre64):

im thinking its 35, but thats just me :)

Parth (parthkohli):

Is this possible? I mean, the number of subsets with 2 elements should ideally be greater than the number of subsets with 3 elements, right? Or did I misinterpret the meaning of the question?

OpenStudy (amistre64):

you might be misreading the question

OpenStudy (amistre64):

think of it like this; what is the number of way of choosing 3 elements from a set of "n", if permuations is not acceptable?

Parth (parthkohli):

\[\binom{n}{3}\]

terenzreignz (terenzreignz):

Fancy ... 35 seems to be correct... But I found no other way to get it except trial and error... There must be a more efficient way :(

OpenStudy (amistre64):

and (n 3) = (n 2) + 14 , if n=?

terenzreignz (terenzreignz):

n = 7

OpenStudy (amistre64):

then (7 4) = 35

hartnn (hartnn):

i too get n=7

Parth (parthkohli):

Ah!

terenzreignz (terenzreignz):

How did you get n = 7 guys? T.T

hartnn (hartnn):

(n choose 2)+14 = (n choose 3)

hartnn (hartnn):

so, yes 7 choose 4 = 35 subsets

OpenStudy (amistre64):

\[\frac{n(n-1)(n-2)}{3!}=\frac{n(n-1)}{2!}+14\] \[n(n-1)(n-2)=3n(n-1)+14(6)\]

terenzreignz (terenzreignz):

oh... that simplifies things :D I must have panicked, @amistre64 It was... a momentary madness~

OpenStudy (amistre64):

or, you could have pascaled a triangle :)

terenzreignz (terenzreignz):

It's a perfect cube?

OpenStudy (amistre64):

not sure if the n stuff above is perfect cubish, but it does have 7 as one of the possible solutions to it; and the only solution that made sense

terenzreignz (terenzreignz):

I'd have just simplified the thing, and then use the rational root test, after all, it's looking for an integer solution. Anyway, I did trial and error'ed it, but that's definitely not the best way to go about it... I learned :D

hartnn (hartnn):

n= 7 is only integer solution, other 2 are complex

OpenStudy (amistre64):

lol, yeah, i just did a wolf run and 7 is the only integer soluton :)

OpenStudy (amistre64):

n(n−1)(n−2)=3n(n−1)+14(6) n^3-3n^2+2n = 3n^2-3n+84 n^3-6n^2+5n - 84 = 0 i think i got to here before going to the pascal

terenzreignz (terenzreignz):

Rational root test... at least it limits your integer solutions :D

OpenStudy (amistre64):

yep, and cardano would be fine as well; let n = (u+6/3) (u+2)^3 - 6(u+2)^2 + 5(u+2) - 84 = 0 u^3 +6u^2 +12u +8 -6u^2 -24u -24 +5u +10 -84 ------------------- u^3 -7u - 90 = 0

OpenStudy (amistre64):

rational roots would at least be simpler :) 1,84 2,42 4,21 3,28 14,6 5,x 7,12

OpenStudy (amistre64):

lol, thatd be more readable as 3,28 6,14 12,7 double the left, half the right

hartnn (hartnn):

hmm...doing manually is quite cumbersome.